I was wondering if anyone can help me out with this problem. "What is the slope of all lines parallel to 4y-2x+1=0?"
Answer by rothauserc(4718)

What is the slope of all lines parallel to 4y-2x+1=0
=====================================================
standard form for a line is y = mx +b
=====================================================
4y-2x+1=0
4y = 2x -1
y = (1/2)x -(1/4)
======================================================
m, the slope of the line is 1/2 which all lines parallel to the given line will have